The first of twelve semiannual payments of $2500 will be made 6.5 years from today. What is the present value of this deferred annuity using a discount rate of 7.5% compounded semiannually? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

  

  Present value $
Answer #1
Step 1 : Value of annuity at the end of year 6 from now
Present Value Of Annuity
= C*[1-(1+i)^-n]/i]
Where,
C= Cash Flow per period
r = Interest rate =7.5%/2 =3.75%
n=number of period
= $2500[ 1-(1+0.0375)^-12 /0.0375]
= $2500[ 1-(1.0375)^-12 /0.0375]
= $2500[ (0.3571) ] /0.0375
= $23,806.73
Step 2 : Value of 23,806.73 today
PV= FV/(1+r)^n
Where,
FV= Future Value
PV = Present Value
r = Interest rate =7.5%/2 =3.75%
n= periods in number =6*2 =12
= $23806.73/( 1+0.0375)^12
=23806.73/1.55545
= $15305.32
Correct Answer = $15305.32